OpenStudy (dragontears1128):

Is this correct? f(x)=5x+5 and g(x)=6x+1 then what is (f+g)(−8)) 5(-8)+5 6(-8)+1 (-35)(-47) =1645

OpenStudy (usukidoll):

(f+g)(x) or f(x) +g(x) is the sum of those two functions f(x) = 5x+5 and g(x) = 6x-1 so f(x) + g(x) would look like this 5x+5+6x-1 combine like terms and we'll get 5x+6x+5-1 11x+4 now we are given that x = -8 so we plug that into the equation 11(-8)+4=-88+4==-84

OpenStudy (usukidoll):

wait I made a typo x( sorry it's almost 4 let me redo this before I go to bed

OpenStudy (usukidoll):

if we had something like (fg)(x) or f(x)g(x) that's multiplication so we will have something like (5x+5)(6x+1) so when x = -8 (5(-8)+5)(6(-8)+1) (-40+5)(-48+1) (-35)(-47)=1645

OpenStudy (usukidoll):

but your question is asking for the sum (f+g)(x) or f(x) +g(x) is the sum of those two functions f(x) = 5x+5 and g(x) = 6x+1 so f(x) + g(x) would look like this 5x+5+6x+1

OpenStudy (usukidoll):

11x+6 so when x = -8 we have 11(-8)+6= -88+6 = -82

OpenStudy (dragontears1128):

thank you